Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Jodrell (Alastair Gunn at Tue Dec 1 08:15:50 MET 1998) Success
Strong RFI. No known problems.

Effelsberg (dgraham at Tue Dec 1 09:49:12 MET 1998) Failure
Lost from begin until 2230 as a thick tape was used. OK with thin tape from 2230 onwards. VCs1-6 have filters for this bandwidth so no power in vc7 and 8.

Torun (K.Borkowski at Tue Dec 1 09:55:45 MET 1998) Success
Early (by about 0.5 minute) stopping of the tape by the low tape sensor at each end of the tape.

Medicina (Maria Rioja at Tue Dec 1 19:24:43 MET 1998) Minor failures
Rain in the beginning of the observations. At Medicina a total of 7 VCs have filters for 0.5 MHz bandwidth. Data registered in VCs 1-7, no data registered from VC 8. Early stopping of the tape, by ca. 1 minute, at the end of each forward and reverse passes.

Westerbork (Tony Foley at Wed Dec 2 11:57:46 MET 1998) Minor failures
We could only observe 3 of the four lines simultaneously, and these were in VCs 1-6. VCs 5,6 and 7 had external filters and I had to increase attenuation to stop these being saturated. Scans after 03:32 were out of hour angle range. The observations used 2 different array setups, one for 1665,1667 and 1612MHz and one for 1665,1667 and 1720MHz.

Onsala (A. Polatidis at Wed Dec 2 14:37:41 MET 1998) Success
No problem with the oobservations. Recorded only on VC1 to VC4 since the other VC are not yet equiped with 0.5 MHz filters

Noto (C.Trigilio at Mon Dec 7 08:25:18 MET 1998) Success

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:47:20 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: ed013.antab. Read the comments in the file to get further information.